OpenStudy (anonymous):

16x − 4x = -48 please show all work

OpenStudy (anonymous):

16x − 4x = -48 12x=-48 12x/12=-48/12 x=-4

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yup